OpenStudy (anonymous):

Solve: 7/(x-1)-5=6/(x^2-1) Please help

OpenStudy (phi):

is this \[ \frac{7}{(x-1)} -5 = \frac{6}{(x^2-1)} \]?

OpenStudy (anonymous):

yeah

OpenStudy (phi):

can you factor (x^2 -1 ) ?

OpenStudy (anonymous):

(x-1)(x+1)

OpenStudy (phi):

\[ \frac{7}{(x-1)} -5 = \frac{6}{(x^2-1)} \] can be written as \[ \frac{7}{(x-1)} -5 = \frac{6}{(x-1)(x+1)} \] I would multiply both sides of the equation (and *all* terms) by (x-1)(x+1) can you do that ?

OpenStudy (anonymous):

yeah 7(x+1)-5(x^2-1)=6 7x+7-5x^2+5=6 -5x^2+7x+2=6

OpenStudy (anonymous):

then do i just factor that?

OpenStudy (phi):

I would first bring the 6 over to the left side -5x^2 + 7 x -4 =0 now (because I don't like leading minus signs), multiply both sides by -1 to get 5x^2 -7x +4 = 0

OpenStudy (anonymous):

ok, so then do i use the quadratic formula?

OpenStudy (phi):

looks like we have to

OpenStudy (phi):

although check your arithmetic 7x+7-5x^2+5=6 this is ok, but the next line looks wrong: -5x^2+7x+2=6

OpenStudy (anonymous):

oh, its +13, isn't it?

OpenStudy (anonymous):

no, +12

OpenStudy (phi):

and then 12 - 6 (bring the 6 over from the right side) I am expecting that whatever we get should factor.

OpenStudy (anonymous):

ok, so then after multiplying by -1 it's 5x^2-7x-6=0

OpenStudy (phi):

yes

OpenStudy (anonymous):

ok, so x=2 and x=-3/5

OpenStudy (phi):

yes, looks good.

OpenStudy (anonymous):

ok, thankyou :)

OpenStudy (phi):

yw